# Disciplined Biconvex Programming

### v1

@@ 1 Introduction
We consider biconvex optimization problems, which consist in optimizing a biconvex objective function subject to a biconvex set constraint. Informally, by saying that a set (or function) is biconvex, we mean that it is convex in each of two blocks of variables when the other block is fixed. Biconvex optimization problems arise in various applications, including machine learning [PT94, UHZB16, ZYHB25], signal and image processing [LWDF09, CVR14], recommender systems [CP09, Joh14], control [SGL94, VB00, JDMV20], and brain computer interfaces [DCP07, SXB14].
@@
Different from convex optimization problems, biconvex optimization problems are in general nonconvex and can be very hard to solve. In fact, many of them have been shown to be NP-hard (see [WYZ12], [TÖ95], and [ZYHB25] for some examples). Nevertheless, various heuristic methods based on alternate convex search (ACS) [DL94] exist for finding good solutions to many biconvex optimization problems in practice. The basic idea of ACS-type methods is to iteratively optimize over one block of variables while keeping the other block fixed, such that each subproblem is convex and can be efficiently solved. Although in the most general case, theoretical convergence of ACS-type methods is only guaranteed to stationary points, these methods turn out to work quite well in practice and hence have become the most popular choice for finding a satisfactory solution to biconvex optimization problems.
@@
Designing and implementing an ACS solver for a specific biconvex optimization problem usually requires significant effort from the user, including properly partitioning the biconvex problem into convex subproblems, implementing robust and efficient solvers for the subproblems, and designing appropriate stopping criteria. Moreover, maintaining numerical stability and ensuring convergence of the ACS procedure typically necessitate integrating appropriate regularizations or modifications into the subproblems to ensure properties such as strong convexity and feasibility. With the help of domain specific languages (DSLs) for convex optimization, such as CVXPY [DB16, AVDB18], specifying and solving the (modified) convex subproblems is largely simplified and automated, while the other steps still require the user to have expert knowledge about convex analysis and programming, and can be tedious and error-prone.
@@
In this paper, we propose a modeling framework for biconvex optimization problems, named disciplined biconvex programming (DBCP), which extends the ideas of disciplined convex programming (DCP) [GBY06] to biconvex problems. DBCP for biconvex optimization is analogous to DCP for convex optimization, which allows users to specify biconvex optimization problems in a natural way based on a small number of syntax rules. When an optimization problem description complies with the DBCP syntax rules, it is guaranteed to be a valid biconvex problem, and more importantly, it can be automatically split and transformed into convex (specifically, DCP-compliant) subproblems and augmented, to which a customized ACS solver is then generated and applied. Like DCP does for convex optimization, DBCP makes it easy to specify and solve biconvex problems, allowing users to quickly prototype and experiment with different biconvex problem formulations, without expertise in (or even knowledge of) convex analysis and solution methods for biconvex problems. In fact, most users might be unaware of the variable partition, transformation, augmentation, and solution processes, from which a solution to the biconvex problem is found. We implement DBCP into a Python package, named dbcp, as an extension of the DSL CVXPY, which is fully open-sourced at
@@ 1.1 Previous and related work
The first notice of biconvexity structure in the context of mathematical programming can be traced back to Falk et al. [FS69] in the 1960s. Only a few papers exist in the literature where biconvex sets are investigated. Most of the corresponding theoretical results can be found in the papers of Aumann and Hart [AH86] and Goh et al. [GTS+94]. Biconvex functions appear regularly in practice and hence have been discussed widely in the literature. Properties of biconvex functions that are most relevant to optimization problems can be found in the works of Goh et al. [GTS+94] and Gorski et al. [GPK07]. In addition, biconvex functions play an important role in martingale theory, and, in particular, in the analysis of Banach and Hilbert spaces [Bur81, Bur86, AH86, Lee93]. Biconvex functions can also be used to derive results on the robust stability of control systems in practical control engineering [GH00a, GH00b]. Thibault [Thi84], Jouak and Thibault [JT85], and Borwein [Bor86] analyzed the continuity and differentiability of measurable biconvex operators in topological vector spaces. Al-Khayyal and Falk [AKF83] published results on the maximization of biconvex functions. In the most general case, very little can be said about the global or even local optimality properties of biconvex optimization problems. Nevertheless, some useful properties of partial optimality conditions (i.e., conditions for stationary points) of biconvex problems are discussed in [WHJ76, GPK07].
@@
Current solution methods for biconvex problems can be roughly categorized into two classes: heuristic methods and global optimization methods. Heuristic methods for biconvex problems aim at finding stationary points of the biconvex objective function, and are mostly based on the idea of alternately optimizing two convex subproblems, i.e., the ACS-type methods. ACS methods are a special case of block relaxation methods [War63, Pow73, DL94] where the variables are divided into disjoint blocks, and in each step, only the variables of an active block are optimized while those of the other blocks are fixed. In particular, for ACS, only two blocks of variables defined by the convex subproblems are activated in cycles. Since the resulting subproblems are convex, efficient convex minimization methods can be used to solve these subproblems. A survey on ACS methods for biconvex optimization problems can be found in [WHJ76]. Gorski et al. [GPK07] showed that under weak assumptions all solution points generated by ACS form a compact connected set and that each of these points is a stationary point of the objective function. However, no better convergence results regarding local or global optimality properties can be obtained in general. Regarding the global optimization methods, Floudas and Visweswaran [FV90] adapted the idea of branch-and-bound [LW66, BM07] to solve biconvex problems with global optimality. Detailed mathematical background and outline of this algorithm are given in [FV90, FV93, Flo95, Flo00], and some basic convergence properties of this method are discussed in [GPK07].
@@
The practical usefulness of biconvex programming in many applications is also increasingly well known. In machine learning, nonnegative matrix factorization [PT94, LS99, UHZB16], dictionary learning [AEB06], and bilinear regression [HKV08, CP09, Joh14] are all well-known biconvex optimization problems, which have been widely applied in representation learning, signal processing, and recommender systems. The very famous $k$-means clustering problem can also be formulated as a biconvex program [Llo82, SMD22, ZYHB25]. Blind deconvolution problems [LWDF09, CVR14] are another important class of biconvex optimization problems for image processing and communication. Other studies such as Mishra et al. [MDC17] and Fosson [Fos18] explored biconvex structures in sparse learning and reweighted regression. Biconvex problems also appear when dealing with bilinear matrix inequalities for bilinear control system synthesis [SGL94, VB00, JDMV20]. Finally, Dyrholm et al. [DCP07] and Shi et al. [SXB14] applied bilinear discriminant component analysis techniques to brain computer interfaces. We discuss some of these applications in more detail in §6.
@@
DCP [GBY06] is a modeling framework for convex optimization problems. If a mathematical optimization problem is written following the DCP syntax ruleset, it is guaranteed to be a valid convex program, and can then be automatically transformed into the standard conic form [AVDB18], which can be handled by generic conic solvers, such as OSQP [SBG+20], ECOS [DCB13], SCS [OCPB16], and Clarabel [GC24]. Many DSLs for convex optimization, such as YALMIP [Lof04], CVX [GB14], Convex.jl [UMZ+14], CVXPY [DB16], and CVXR [FNB20], are based on the DCP ruleset. Over the years, DCP has been extended to handle stochastic [AKDB15], convex-concave [SDGB16], geometric [ADB19], quasiconvex [AB20], and saddle [SLB24] problems. The most related extension of DCP to our work is disciplined multi-convex programming [SDU+17], which provides a DSL for specifying and solving multi-convex optimization problems via block relaxation methods. However, it seems to be no longer actively supported or maintained. In this paper, we work with CVXPY and focus on biconvex problems only, since most multi-convex problems in practice can be formulated as biconvex problems by grouping variables appropriately.
@@ 2 Biconvex programming
In this section, we provide formal definitions for biconvex sets, biconvex functions, and biconvex optimization problems, to establish our notation, and briefly review some of the basic properties of these objects. To start with, let ${\cal X{}}\subseteq{\bf R{}}^{n}$ and ${\cal Y{}}\subseteq{\bf R{}}^{k}$ be two nonempty closed convex sets. We should also note that throughout this paper, we use the convention that functions are implicitly extended-valued, i.e., by writing $f\colon A\to{\bf R{}}$, we really mean $f\colon A\to{\bf R{}}\cup\{\infty\}$. In other words, a function defined on $A$ can take the value $\infty$ for some points in $A$ [Roc70, BV04].
@@ 2.1 Biconvex sets
An important property of biconvex sets is that, similar to convex sets, the intersection of an arbitrary collection of biconvex sets is still a biconvex set [AH86, GPK07].
@@ 2.2 Biconvex functions
Now we list some basic properties of biconvex functions. We do not provide formal proofs here; interested readers may refer to [GTS+94] and [GPK07] for the corresponding proofs and discussion of the following statements (which are more or less obvious with the basic convex analysis toolbox [Roc70]). Similar to convex functions, the sublevel sets of a biconvex function $f\colon{\cal X{}}\times{\cal Y{}}\to{\bf R{}}$, given by
@@
  • Composition. If $h\colon{\bf R{}}\to{\bf R{}}$ is convex and nondecreasing, and $g\colon{\cal X{}}\times{\cal Y{}}\to{\bf R{}}$ is biconvex, then the function $f\colon{\cal X{}}\times{\cal Y{}}\to{\bf R{}}$, given by $f(x,y)=h(g(x,y))$ is biconvex. This property can be easily extended to multivariate functions $h\colon{\bf R{}}^{m}\to{\bf R{}}$ that are convex and nondecreasing in each argument. Interested readers may refer to [BV04, §3.2.4] for more details.
@@ 2.3 Biconvex optimization problems
It can be shown that for a differentiable biconvex optimization problem, every stationary point of $f_{0}$ over $\cal D{}$} is partially optimal, and vise versa [GPK07]. However, such a point is not necessarily a local optimum, as stationary points can be saddle points of the objective function. Some necessary conditions for a partially optimal point of a biconvex problem being a local optimum are discussed in [GPK07], but in general, no stronger results can be obtained. Albeit a weak notion of optimality, partially optimal points for biconvex problems is still widely used in practice, and turns out to be good enough for many applications.
@@ 3.1 Alternate convex search
ACS is a minimization method as a special case of block relaxation methods [DL94], where the variables are divided into two disjoint blocks, and in each step, only the variables of an active block are optimized while those of the other block are fixed. Specifically, for a biconvex optimization problem of the form (2.1), ACS iterates between solving the following two convex subproblems:
@@
Gorski et al. [GPK07] showed that under weak assumptions, the sequence of objective function values ${\{f_{0}(x^{(k)},y^{(k)})\}}_{k=0}^{\infty}$ generated by ACS is monotonically nonincreasing and convergent. Furthermore, if the sequence ${\{(x^{(k)},y^{(k)})\}}_{k=0}^{\infty}$ converges to $(x^{\star},y^{\star})$, then $(x^{\star},y^{\star})$ is a stationary point of $f_{0}$ (and hence, is a partially optimal point of (2.1)). We should note that ACS is sensitive to initialization, i.e., with different initial points $(x^{(0)},y^{(0)})$, the sequence ${\{(x^{(k)},y^{(k)})\}}_{k=0}^{\infty}$ generated through algorithm 3.1 may converge to different stationary points of $f_{0}$ with (probably) different function values.
@@ 3.2 Proximal regularizations
where $\lambda\geq 0$ is the regularization parameter. When $\lambda=0$, the updates (3.7) and (3.8) reduce to (3.3) and (3.4), respectively. When $\lambda>0$, the proximal terms in (3.7) and (3.8) can be interpreted as adding soft trust region constraints to the respective optimization variables, which penalize large changes of the variables between two consecutive iterations. By choosing sufficiently large $\lambda$, the corresponding optimization problems in (3.7) and (3.8) become strongly convex, which can help improve numerical stability in practice [BPC+11, PB14]. It is also observed that adding the additional proximal regularizers can sometimes lead to better final points, i.e., points with lower objective values [SDU+17], compared to those from the original ACS procedure.
@@ 3.3 Initialization
with variables $x\in{\cal X{}}$ and $y\in{\cal Y{}}$. Solving (3.9) directly can be as hard as solving the original biconvex problem (2.1), and, actually, it can even be NP-hard [TÖ95]. Here, we consider the following heuristic via relaxation to find a feasible starting point:

### v2

@@ 1 Introduction
We consider biconvex optimization problems, which consist in optimizing a biconvex objective function subject to a biconvex set constraint. Informally, by saying that a set (or function) is biconvex, we mean that it is convex in each of two blocks of variables when the other block is fixed. Biconvex optimization problems arise in various applications, including machine learning [paatero1994positive, udell2016generalized, zhu2025multi], signal and image processing [levin2009understanding, chaudhuri2014blind], recommender systems [chu2009personalized, johnson2014logistic], control [safonov1994control, vanantwerp2000tutorial, javanmardi2020bilinear], and brain computer interfaces [dyrholm2007bilinear, shi2014sparse].
@@
Different from convex optimization problems, biconvex optimization problems are in general nonconvex and can be very hard to solve. In fact, many of them have been shown to be NP-hard (see [wen2012solving], [toker1995np], and [zhu2025multi] for some examples). Nevertheless, various heuristic methods based on alternate convex search (ACS) [de1994block] exist for finding good solutions to many biconvex optimization problems in practice. The basic idea of ACS-type methods is to iteratively optimize over one block of variables while keeping the other block fixed, such that each subproblem is convex and can be efficiently solved. Although in the most general case, theoretical convergence of ACS-type methods is only guaranteed to stationary points, these methods turn out to work quite well in practice and hence have become the most popular choice for finding a satisfactory solution to biconvex optimization problems.
@@
Designing and implementing an ACS solver for a specific biconvex optimization problem usually requires significant effort from the user, including properly partitioning the biconvex problem into convex subproblems, implementing robust and efficient solvers for the subproblems, and designing appropriate stopping criteria. Moreover, maintaining numerical stability and ensuring convergence of the ACS procedure typically necessitate integrating appropriate regularizations or modifications into the subproblems to ensure properties such as strong convexity and feasibility. With the help of domain specific languages (DSLs) for convex optimization, such as CVXPY [diamond2016cvxpy, agrawal2018rewriting], specifying and solving the (modified) convex subproblems is largely simplified and automated, while the other steps still require the user to have expert knowledge about convex analysis and programming, and can be tedious and error-prone.
@@
In this paper, we propose a modeling framework for biconvex optimization problems, named disciplined biconvex programming (DBCP), which extends the ideas of disciplined convex programming (DCP) [grant2006disciplined] to biconvex problems. DBCP for biconvex optimization is analogous to DCP for convex optimization, which allows users to specify biconvex optimization problems in a natural way based on a small number of syntax rules. When an optimization problem description complies with the DBCP syntax rules, it is guaranteed to be a valid biconvex problem, and more importantly, it can be automatically split and transformed into convex (specifically, DCP-compliant) subproblems and augmented, to which a customized ACS solver is then generated and applied. Like DCP does for convex optimization, DBCP makes it easy to specify and solve biconvex problems, allowing users to quickly prototype and experiment with different biconvex problem formulations, without expertise in (or even knowledge of) convex analysis and solution methods for biconvex problems. In fact, most users might be unaware of the variable partition, transformation, augmentation, and solution processes, from which a solution to the biconvex problem is found. We implement DBCP into a Python package, named dbcp, as an extension of the DSL CVXPY, which is fully open-sourced at
@@ 1.1 Previous and related work
The first notice of biconvexity structure in the context of mathematical programming can be traced back to Falk et al. [falk1969algorithm] in the 1960s. Only a few papers exist in the literature where biconvex sets are investigated. Most of the corresponding theoretical results can be found in the papers of Aumann and Hart [aumann1986bi] and Goh et al. [goh1994biaffine]. Biconvex functions appear regularly in practice and hence have been discussed widely in the literature. Properties of biconvex functions that are most relevant to optimization problems can be found in the works of Goh et al. [goh1994biaffine] and Gorski et al. [gorski2007biconvex]. In addition, biconvex functions play an important role in martingale theory, and, in particular, in the analysis of Banach and Hilbert spaces [burkholder1981geometrical, burkholder1986martingales, aumann1986bi, lee1993burkholder]. Biconvex functions can also be used to derive results on the robust stability of control systems in practical control engineering [geng2000robusta, geng2000robustb]. Thibault [thibault1984continuity], Jouak and Thibault [jouak1985directional], and Borwein [borwein1986partially] analyzed the continuity and differentiability of measurable biconvex operators in topological vector spaces. Al-Khayyal and Falk [al1983jointly] published results on the maximization of biconvex functions. In the most general case, very little can be said about the global or even local optimality properties of biconvex optimization problems. Nevertheless, some useful properties of partial optimality conditions (i.e., conditions for stationary points) of biconvex problems are discussed in [wendell1976minimization, gorski2007biconvex].
@@
Current solution methods for biconvex problems can be roughly categorized into two classes: heuristic methods and global optimization methods. Heuristic methods for biconvex problems aim at finding stationary points of the biconvex objective function, and are mostly based on the idea of alternately optimizing two convex subproblems, i.e., the ACS-type methods. ACS methods are a special case of block relaxation methods [warga1963minimizing, powell1973search, de1994block] where the variables are divided into disjoint blocks, and in each step, only the variables of an active block are optimized while those of the other blocks are fixed. In particular, for ACS, only two blocks of variables defined by the convex subproblems are activated in cycles. Since the resulting subproblems are convex, efficient convex minimization methods can be used to solve these subproblems. A survey on ACS methods for biconvex optimization problems can be found in [wendell1976minimization]. Gorski et al. [gorski2007biconvex] showed that under weak assumptions all solution points generated by ACS form a compact connected set and that each of these points is a stationary point of the objective function. However, no better convergence results regarding local or global optimality properties can be obtained in general. Regarding the global optimization methods, Floudas and Visweswaran [floudas1990global] adapted the idea of branch-and-bound [lawler1966branch, boyd2007branch] to solve biconvex problems with global optimality. Detailed mathematical background and outline of this algorithm are given in [floudas1990global, floudas1993primal, floudas1995nonlinear, floudas2000deterministic], and some basic convergence properties of this method are discussed in [gorski2007biconvex].
@@
The practical usefulness of biconvex programming in many applications is also increasingly well known. In machine learning, nonnegative matrix factorization [paatero1994positive, lee1999learning, udell2016generalized], dictionary learning [aharon2006k], and bilinear regression [hu2008collaborative, chu2009personalized, johnson2014logistic] are all well-known biconvex optimization problems, which have been widely applied in representation learning, signal processing, and recommender systems. The very famous $k$-means clustering problem can also be formulated as a biconvex program [lloyd1982least, so2022convergence, zhu2025multi]. Blind deconvolution problems [levin2009understanding, chaudhuri2014blind] are another important class of biconvex optimization problems for image processing and communication. Other studies such as Mishra et al. [mishra2017sequential] and Fosson [fosson2018biconvex] explored biconvex structures in sparse learning and reweighted regression. Biconvex problems also appear when dealing with bilinear matrix inequalities for bilinear control system synthesis [safonov1994control, vanantwerp2000tutorial, javanmardi2020bilinear]. Finally, Dyrholm et al. [dyrholm2007bilinear] and Shi et al. [shi2014sparse] applied bilinear discriminant component analysis techniques to brain computer interfaces. We discuss some of these applications in more detail in §6.
@@
DCP [grant2006disciplined] is a modeling framework for convex optimization problems. If a mathematical optimization problem is written following the DCP syntax ruleset, it is guaranteed to be a valid convex program, and can then be automatically transformed into the standard conic form [agrawal2018rewriting], which can be handled by generic conic solvers, such as OSQP [stellato2020osqp], ECOS [domahidi2013ecos], SCS [o2016conic], and Clarabel [goulart2024clarabel]. Many DSLs for convex optimization, such as YALMIP [lofberg2004yalmip], CVX [grant2014cvx], Convex.jl [udell2014convex], CVXPY [diamond2016cvxpy], and CVXR [fu2020cvxr], are based on the DCP ruleset. Over the years, DCP has been extended to handle stochastic [ali2015disciplined], convex-concave [shen2016disciplined], geometric [agrawal2019disciplined], quasiconvex [agrawal2020disciplined], and saddle [schiele2024disciplined] problems. The most related extension of DCP to our work is disciplined multi-convex programming [shen2017disciplined], which provides a DSL for specifying and solving multi-convex optimization problems via block relaxation methods. However, it seems to be no longer actively supported or maintained. In this paper, we work with CVXPY and focus on biconvex problems only, since most multi-convex problems in practice can be formulated as biconvex problems by grouping variables appropriately.
@@ 2 Biconvex programming
In this section, we provide formal definitions for biconvex sets, biconvex functions, and biconvex optimization problems, to establish our notation, and briefly review some of the basic properties of these objects. To start with, let ${\cal X{}}\subseteq{\bf R{}}^{n}$ and ${\cal Y{}}\subseteq{\bf R{}}^{k}$ be two nonempty closed convex sets. We should also note that throughout this paper, we use the convention that functions are implicitly extended-valued, i.e., by writing $f\colon A\to{\bf R{}}$, we really mean $f\colon A\to{\bf R{}}\cup\{\infty\}$. In other words, a function defined on $A$ can take the value $\infty$ for some points in $A$ [rockafellar1970convex, boyd2004convex].
@@ 2.1 Biconvex sets
An important property of biconvex sets is that, similar to convex sets, the intersection of an arbitrary collection of biconvex sets is still a biconvex set [aumann1986bi, gorski2007biconvex].
@@ 2.2 Biconvex functions
Now we list some basic properties of biconvex functions. We do not provide formal proofs here; interested readers may refer to [goh1994biaffine] and [gorski2007biconvex] for the corresponding proofs and discussion of the following statements (which are more or less obvious with the basic convex analysis toolbox [rockafellar1970convex]). Similar to convex functions, the sublevel sets of a biconvex function $f\colon{\cal X{}}\times{\cal Y{}}\to{\bf R{}}$, given by
@@
  • Composition. If $h\colon{\bf R{}}\to{\bf R{}}$ is convex and nondecreasing, and $g\colon{\cal X{}}\times{\cal Y{}}\to{\bf R{}}$ is biconvex, then the function $f\colon{\cal X{}}\times{\cal Y{}}\to{\bf R{}}$, given by $f(x,y)=h(g(x,y))$ is biconvex. This property can be easily extended to multivariate functions $h\colon{\bf R{}}^{m}\to{\bf R{}}$ that are convex and nondecreasing in each argument. Interested readers may refer to [boyd2004convex, §3.2.4] for more details.
@@ 2.3 Biconvex optimization problems
It can be shown that for a differentiable biconvex optimization problem, every stationary point of $f_{0}$ over $\cal D{}$} is partially optimal, and vice versa [gorski2007biconvex]. However, such a point is not necessarily a local optimum, as stationary points can be saddle points of the objective function. Some necessary conditions for a partially optimal point of a biconvex problem being a local optimum are discussed in [gorski2007biconvex], but in general, no stronger results can be obtained. Albeit a weak notion of optimality, partially optimal points for biconvex problems is still widely used in practice, and turns out to be good enough for many applications.
@@ 3.1 Alternate convex search
ACS is a minimization method as a special case of block relaxation methods [de1994block], where the variables are divided into two disjoint blocks, and in each step, only the variables of an active block are optimized while those of the other block are fixed. Specifically, for a biconvex optimization problem of the form (2.1), ACS iterates between solving the following two convex subproblems:
@@
Gorski et al. [gorski2007biconvex] showed that under weak assumptions, the sequence of objective function values ${\{f_{0}(x^{(k)},y^{(k)})\}}_{k=0}^{\infty}$ generated by ACS is monotonically nonincreasing and convergent. Furthermore, if the sequence ${\{(x^{(k)},y^{(k)})\}}_{k=0}^{\infty}$ converges to $(x^{\star},y^{\star})$, then $(x^{\star},y^{\star})$ is a stationary point of $f_{0}$ (and hence, is a partially optimal point of (2.1)). We should note that ACS is sensitive to initialization, i.e., with different initial points $(x^{(0)},y^{(0)})$, the sequence ${\{(x^{(k)},y^{(k)})\}}_{k=0}^{\infty}$ generated through algorithm 3.1 may converge to different stationary points of $f_{0}$ with (probably) different function values.
@@ 3.2 Proximal regularizations
where $\lambda\geq 0$ is the regularization parameter. When $\lambda=0$, the updates (3.7) and (3.8) reduce to (3.3) and (3.4), respectively. When $\lambda>0$, the proximal terms in (3.7) and (3.8) can be interpreted as adding soft trust region constraints to the respective optimization variables, which penalize large changes of the variables between two consecutive iterations. By choosing sufficiently large $\lambda$, the corresponding optimization problems in (3.7) and (3.8) become strongly convex, which can help improve numerical stability in practice [boyd2011distributed, parikh2014proximal]. It is also observed that adding the additional proximal regularizers can sometimes lead to better final points, i.e., points with lower objective values [shen2017disciplined], compared to those from the original ACS procedure.
@@ 3.3 Initialization
with variables $x\in{\cal X{}}$ and $y\in{\cal Y{}}$. Solving (3.9) directly can be as hard as solving the original biconvex problem (2.1), and, actually, it can even be NP-hard [toker1995np]. Here, we consider the following heuristic via relaxation to find a feasible starting point: